OpenStudy (anonymous):

perform and simplify completely 2x 6x+3 ----- * ------- 4x+2 5

OpenStudy (anonymous):

I think you will find it easier if you realize that \[6x + 3 = 3(2x+1)\] and \[4x+2 = 2(2x+1)\]

OpenStudy (anonymous):

that is a little better but im still lost with all of this

OpenStudy (anonymous):

Cross cancel?

OpenStudy (anonymous):

like i said....all seems like a foreign language to me....do you mean to use the "foil" method?

OpenStudy (anonymous):

No, I mean cross cancel. Like when you have: \[\frac{2}{3} \times \frac{3}{4}\]

OpenStudy (anonymous):

Or if you aren't familiar with that then at least you can cancel common factors

OpenStudy (anonymous):

\[\frac{2x}{4x+2} \times \frac{6x+3}{5}\]\[=\frac{2x}{2(2x+1)} \times \frac{3(2x+1)}{5}\]\[= {3(2)(x)(2x+1) \over 2(5)(2x+1)}\] The 2x + 1 factor cancels from the top and bottom, and the twos cancel from the top and bottom.